51 eV of energy is supplied to He+ in ground state. Find out the total possible no. of spectral lines observed if transition takes place to ground state.













































(1) 0
(2) 1
(3) 2
(4) 3
(5) 4
(6) 5
(7) 6
(8) 7
(9) 8
(10) 9

Grade:12

1 Answers

saloni sharma
10 years ago
Since the energy supplied -51evtherefore, 51= 13.6•4 - 13.6•4/n^251/13.6•4= 1-1/n^2.9375 = 1- 1/n^21/n^2 = 1- 0.9375n^2= 1/0.0625n = 4no. Of spectral lines = (n2 -n1)(n2-n1 + 1)/2 = 6